# Stronger, Steadier & Superior: Geometric Consistency in Depth VFM Forges Domain Generalized Semantic Segmentation

### v2

@@ Abstract
Vision Foundation Models (VFMs) have delivered remarkable performance in Domain Generalized Semantic Segmentation (DGSS). However, recent methods often overlook the fact that visual cues are susceptible, whereas the underlying geometry remains stable, rendering depth information more robust. In this paper, we investigate the potential of integrating depth information with features from VFMs, to improve the geometric consistency within an image and boost the generalization performance of VFMs. We propose a novel fine-tuning DGSS framework, named DepthForge, which integrates the visual cues from frozen DINOv2 or EVA02 and depth cues from frozen Depth Anything V2. In each layer of the VFMs, we incorporate depth-aware learnable tokens to continuously decouple domain-invariant visual and spatial information, thereby enhancing depth awareness and attention of the VFMs. Finally, we develop a depth refinement decoder and integrate it into the model architecture to adaptively refine multi-layer VFM features and depth-aware learnable tokens. Extensive experiments are conducted based on various DGSS settings and five different datsets as unseen target domains. The qualitative and quantitative results demonstrate that our method significantly outperforms alternative approaches with stronger performance, steadier visual-spatial attention, and superior generalization ability. In particular, DepthForge exhibits outstanding performance under extreme conditions (e.g., night and snow). Code is available at https://github.com/anonymouse-xzrptkvyqc/DepthForge.
@@ 1 Introduction
In this paper, we introduce DepthForge, a robust fine-tuning and optimization strategy for domain generalization semantic segmentation, as shown in Fig. 1. Specifically, we employ two frozen VFMs: Depth Anything V2[49, 48] is dedicated to extract depth information, and DINOv2[29] or EVA02[14] are used to process RGB visual features. We enhance the generalization of the visual features at each layer of the VFMs by integrating the corresponding depth cues. More importantly, we introduce depth-aware learnable tokens that enable the model to learn spatial structure invariance, thereby ensuring consistent performance across different domains. The depth and visual features extracted from frozen VFMs are used to optimize the representation of our learnable tokens, thus producing more robust visual-spatial attention features than those based solely on visual cues and forging geometric consistency, as shown in Fig. 1(b). However, we observe that the frozen VFMs only provide static features, they are incapable of refining the feature maps during training. This limitation leads to persistent errors in the learnable tokens, causing the optimization process to drift in incorrect directions that are difficult to rectify. To this end, we propose a depth refinement decoder that dynamically adjusts the multi-scale features, combining prior features and dynamic features to establish high-quality paired feature relationships at different depth spaces.
@@ 2.1 Domain Generalized Semantic Segmentation
Figure 3: DepthForge framework. It incoprates the features from frozen VFM and depth VFM into our Depth-aware Learnable Tokens to generate enhanced features $f_{i+1}=f^{v}_{i}+DepthForge(f^{v}_{i},f^{d}_{i})$ for each layer. In our tunable DepthForge, We first design a Depth Awareness module that intergrates visual and depth cues into learnable tokens. Next, we propose an Attention Optimization module to strengthen the discriminability of instances across different spatial locations. Finally, multi-scale features are fused in Depth refinement Decoder to adapt to various scales and spatial connexts. Our DepthForge establishes direct connections to different instances in real-world locations, facilitating both feature refinement and improved discriminability.
@@ 2.2 RGB-D Segmentation
Combining RGB with depth information has shown substantial benefits across various visual tasks [16, 8, 6, 38]. In object detection, RGB-D effectively mitigates limitations inherent to RGB data under complex scenes by providing robust geometric and spatial cues, thereby enhancing localization accuracy and model robustness [42, 58]. In semantic segmentation, depth information assists models in capturing spatial structures and geometric contours of objects, alleviating issues related to imprecise object boundary delineation. Particularly in scenarios with blurred boundaries, insufficient illumination, or texture scarcity, the incorporation of depth information substantially improves segmentation accuracy and stability [47, 59, 53, 13]. Recent studies have further indicated that even in domain generalization tasks, integrating RGB and depth features consistently improves the model’s generalization performance on unseen domains [18, 37, 26, 54]. Therefore, exploring the integration of depth information to address the limitations of RGB-only approaches represents a promising and essential research direction within DGSS.
@@ 3 Preliminary
We use the frozen features $f^{v}_{i}$ from VFM for visual embedding, and the frozen festures $f^{d}_{i}$ from depth VFM for depth embedding. A straightforward thought might simply add $f^{d}_{i}$ to $f^{v}_{i}$, or incorporate $f^{d}_{i}$ into the leanable token $T_{i}$, However, these methods lack flexibility and introduce significant inter-layer feature differences that prevent convergence, as shown in Fig. 2. In contrast, we propose DepthForge to adaptively fuse visual and depth cues, denoted as:
@@
where $W^{v}_{i}$ and $W^{d}_{i}$ represent the weight matrix from VFMs and depth VFM, respectively. $\varepsilon$ denote the adaptation of the visual and depth signals. This approach allows us to more effectively utilize the powerful capacities of multi-modal VFMs to forge robust refined features and instance discriminability.
@@ 4 DepthForge
This paper introduces the DepthForge strategy, a novel approach for domain generalized semantic segmenation. As shown in Fig. 3, DepthForge consists of three main parts: (1) Depth Awareness; (2) Attention Optimization; (3) Depth Refinement Decoder. We detail the designs as follows.
@@ 4.1 Depth Awareness
We employ a set of learnable tokens $T=\{T_{i}\in\mathbb{R}^{m\times c}|i\in[1,N]\}$ to adaptively learn scene knowledge. As we consider, relying solely on visual cues can be limiting and prone to failure. Therefore, we utilize visual and depth cues to construct learnable tokens, which we term depth-aware learnable tokens. We integrate $T_{i}$ with $f^{v}_{i}$ and $f^{d}_{i}$ using an attention-based mechanism, prompting tailored adjustments to bridge the gap between source and target domains.
We use $f^{v,d}_{i}$ as the $query$, and $T_{i}$ as the $key$ and the $value$ A dot-product operation is applied to generate awarness map $A^{v,d}_{i}$, and the $Softmax$ function is employed to align each patch with a unique instance:
$A^{v,d}_{i}=Softmax(\frac{f^{v,d}_{i}\times T^{\top}_{i}}{\sqrt{c}}).$ (5)
The $A^{v,d}_{i}$ provides insights into both the visual cues and depth retrieval, qualitatively evaluating the relationship between various tokens and visual or depth cues, respectively. Subsequently, we incorporate depth awareness bias $A^{d}_{i}$ into the visual features $A^{v}_{i}$ to provide additional spatial relationships, thereby establishing a latent connection between pixels and the real world. The process is mathematically formulated as $A_{i}=A^{v}_{i}+A^{d}_{i}$.
@@ 4.3 Depth Refinement Decoder
Previous methods rely solely on the output of the final layer $f_{N}$ for semantic segmentation predictions. In contrast, we fuse the outputs from layers $L_{1}-L_{N}$ to refine the depth features. Specifically, we process these multi-layer features with an MLP with two differnet fully-connected layers $\phi$ and $\varphi$ and then concatenate them after a RelU activation function:
@@ 5.4 Ablation Studies
Depth awareness and attention optimization are significant for depth-aware learnable tokens. We find that simply incorporating a spatial cue bias into the original learnable tokens ($+f^{d}_{i}$) makes almost no difference. However, when we transform the original learnable tokens into depth-aware learnable tokens using the DepthForge framework, we achieve an approximate 2% improvement in the target domains ACDC and Mapillary, as shown in Tab. 4. The depth-aware learnable tokens compensate for the shortcomings of visual cues and establish more robust visual-spatial attention. This conclusion is supported by qualitative visualizations in Fig. 7.
@@ 6 Conclusion
Limitations. The proposed method increases the parameter count by approximately 4M compared to previous work, but DepthForge remains parameter-efficient to balance performance and efficiency. However, the incorporation of depth cues may cause interference between the two VFMs, and in the future, we need to consider a unified architecture to achieve an implicit and efficient fusion.

### v3

@@ Abstract
Vision Foundation Models (VFMs) have delivered remarkable performance in Domain Generalized Semantic Segmentation (DGSS). However, recent methods often overlook the fact that visual cues are susceptible, whereas the underlying geometry remains stable, rendering depth information more robust. In this paper, we investigate the potential of integrating depth information with features from VFMs, to improve the geometric consistency within an image and boost the generalization performance of VFMs. We propose a novel fine-tuning DGSS framework, named DepthForge, which integrates the visual cues from frozen DINOv2 or EVA02 and depth cues from frozen Depth Anything V2. In each layer of the VFMs, we incorporate depth-aware learnable tokens to continuously decouple domain-invariant visual and spatial information, thereby enhancing depth awareness and attention of the VFMs. Finally, we develop a depth refinement decoder and integrate it into the model architecture to adaptively refine multi-layer VFM features and depth-aware learnable tokens. Extensive experiments are conducted based on various DGSS settings and five different datasets as unseen target domains. The qualitative and quantitative results demonstrate that our method significantly outperforms alternative approaches with stronger performance, steadier visual-spatial attention, and superior generalization ability. In particular, DepthForge exhibits outstanding performance under extreme conditions (e.g., night and snow). Code is available at https://github.com/SY-Ch/DepthForge.
@@ 1 Introduction
In this paper, we introduce DepthForge, a robust fine-tune and optimization strategy for domain generalized semantic segmentation, as shown in Fig. 1. Specifically, we employ two frozen VFMs: Depth Anything V2[50, 49] is dedicated to extract depth information, and DINOv2[30] or EVA02[15] are used to process RGB visual features. We enhance the generalization of the visual features at each layer of the VFMs by integrating the corresponding depth cues. More importantly, we introduce depth-aware learnable tokens that enables the model to learn spatial structure invariance, thereby ensuring consistent performance across different domains. The depth and visual features extracted from frozen VFMs are used to optimize the representation of our learnable tokens, thus producing more robust visual-spatial attention features than those based solely on visual cues and forging geometric consistency, as shown in Fig. 1(b). However, we observe the frozen VFMs only provide static features, they are incapable of refining the feature maps during training. This limitation leads to persistent errors in the learnable tokens, causing the optimization process to drift in incorrect directions that are difficult to rectify. To this end, we propose a depth refinement decoder that dynamically adjusts the multi-scale features, combining prior and dynamic features to establish high-quality paired feature relationships at different depth spaces.
@@ 2.1 Domain Generalized Semantic Segmentation
Figure 3: DepthForge framework. It incorporates the features from frozen VFM and depth VFM into our Depth-aware Learnable Tokens to generate enhanced features $f_{i+1}=f^{v}_{i}+DepthForge(f^{v}_{i},f^{d}_{i})$ for each layer. In our tunable DepthForge, We first design a Depth Awareness module that integrates visual and depth cues into learnable tokens. Next, we propose an Attention Optimization module to strengthen the discriminability of instances across different spatial locations. Finally, multi-scale features are fused in Depth Refinement Decoder to adapt to various scales and spatial connections. Our DepthForge establishes direct connections to different instances in real-world locations, facilitating both feature refinement and improved discriminability.
@@ 2.2 RGB-D Segmentation
Combining RGB with depth information has shown substantial benefits across various visual tasks [17, 9, 6, 39, 8]. In object detection, RGB-D effectively mitigates limitations inherent to RGB data under complex scenes by providing robust geometric and spatial cues, thereby enhancing localization accuracy and model robustness [43, 59]. In semantic segmentation, depth information assists models in capturing spatial structures and geometric contours of objects, alleviating issues related to imprecise object boundary delineation. Particularly in scenarios with blurred boundaries, insufficient illumination, or texture scarcity, the incorporation of depth information substantially improves segmentation accuracy and stability [48, 60, 54, 14]. Recent studies have further indicated that even in domain generalization tasks, integrating RGB and depth features consistently improves the model’s generalization performance on unseen domains [19, 38, 27, 55]. Therefore, exploring the integration of depth information to address the limitations of RGB-only approaches represents a promising and essential research direction within DGSS.
@@ 3 Preliminary
We use the frozen features $f^{v}_{i}$ from VFM for visual embedding, and the frozen features $f^{d}_{i}$ from depth VFM for depth embedding. A straightforward thought might simply add $f^{d}_{i}$ to $f^{v}_{i}$, or incorporate $f^{d}_{i}$ into the leanable token $T_{i}$. However, these methods lack flexibility and introduce significant inter-layer feature differences that prevent convergence, as shown in Fig. 2. In contrast, we propose DepthForge to adaptively fuse visual and depth cues, denoted as:
@@
where $W^{v}_{i}$ and $W^{d}_{i}$ represent the weight matrix from VFMs and depth VFM, respectively. $\varepsilon$ denotes the adaptation of the visual and depth signals. This approach allows us to more effectively utilize the powerful capacities of multi-modal VFMs to forge robust refined features and instance discriminability.
@@ 4 DepthForge
This paper introduces the DepthForge strategy, a novel approach for domain generalized semantic segmentation. As shown in Fig. 3, DepthForge consists of three main parts: (1) Depth Awareness; (2) Attention Optimization; (3) Depth Refinement Decoder. We detail the designs as follows.
@@ 4.1 Depth Awareness
We employ a set of learnable tokens $T=\{T_{i}\in\mathbb{R}^{m\times c}|i\in[1,N]\}$ to adaptively learn scene knowledge. As we consider, relying solely on visual cues can be limiting and prone to failure. Therefore, we utilize visual and depth cues to construct learnable tokens, which we term depth-aware learnable tokens. We integrate $T_{i}$ with $f^{v}_{i}$ and $f^{d}_{i}$ using an attention-based mechanism, prompting tailored adjustments to bridge the gap between different domains.
We use $f^{v,d}_{i}$ as the $query$, and $T_{i}$ as the $key$ and the $value$ A dot-product operation is applied to generate awareness map $A^{v,d}_{i}$, and the $Softmax$ function ($S$) is employed to align each patch with a unique instance. The $A^{v,d}_{i}$ provides insights into both the visual cues and depth retrieval, qualitatively evaluating the relationship between various tokens and visual or depth cues, respectively. Subsequently, we incorporate depth awareness bias $A^{d}_{i}$ into the visual features $A^{v}_{i}$ to provide additional spatial relationships, thereby establishing a latent connection between pixels and the real world. The process is mathematically formulated as:
$A_{i}=S(\frac{f^{v}_{i}\times(T^{v})^{\top}_{i}}{\sqrt{c}})+\lambda S(\frac{f^{d}_{i}\times(T^{d})^{\top}_{i}}{\sqrt{c}}).$ (5)
@@ 4.3 Depth Refinement Decoder
Previous methods rely solely on the output of the final layer $f_{N}$ for semantic segmentation predictions. In contrast, we fuse the outputs from layers $L_{1}-L_{N}$ to refine the depth features. Specifically, we process these multi-layer features with an MLP with two different fully-connected layers $\phi$ and $\varphi$ and then concatenate them after a ReLU activation function:
@@ 5.3 Comparison with State-of-The-Art Methods
No. Design Avg. Gain (%) Rain Night Snow Rog
1 No Depth - 70.6 55.9 79.5 72.5
2 Concatenation -1.1 69.8 55.0 78.3 71.2
3 Depth Token 1.6 72.3 57.8 80.7 74.0
4 Config1 0.2 70.8 56.1 79.4 72.8
5 Config2 0.7 71.5 56.8 80.1 73.0
6 Config3, Ours 3.5 75.4 60.4 81.2 75.4
7 w/o Scale Factor 2.8 74.8 59.3 80.9 74.8
Table 5: Performance comparison of the different configurations. Top three results are highlighted as best, second, and third, respectively. (%)
@@ 5.4 Ablation Studies
Depth awareness and attention optimization are significant for depth-aware learnable tokens. We find that simply incorporating a spatial cue bias into the original learnable tokens ($+f^{d}_{i}$) makes almost no difference. However, when we transform the original learnable tokens into depth-aware learnable tokens using the DepthForge framework, we achieve an approximate 2% improvement in the target domains ACDC and Mapillary, as shown in Tab. 4. The depth-aware learnable tokens compensate for the shortcomings of visual cues and establish more robust visual-spatial attention. Compared to different configurations, our design allows gradients to be backpropagated through the tokens alone, enables each modality to contribute independently based on its own feature representations, preventing dominance from one during optimization, as shown in Tab. 5 and Fig. 2. This conclusion is supported by qualitative visualizations in Fig. 7.
@@ 6 Conclusion
Acknowledgements
This work was supported in part by the Natural Science Foundation of Xiamen, China, under Grant 3502Z202373036; in part by the National Natural Science Foundation of China under Grant 42371457; in part by the Key Project of Natural Science Foundation of Fujian Province, China, under Grant 2022J02045.